OpenStudy (anonymous):

Janet Home went to Citizen Bank. She borrowed $7,000 at a rate of 8 percent. The date of the loan was September 20. Janet hoped to repay the loan on January 20. Assuming the loan is based on ordinary interest, Janet will pay back interest on January 20:

OpenStudy (anonymous):

I assume 8% interest is per year. Is the problem asking to find the interest paid on January 20th?

OpenStudy (anonymous):

Interest per month = 8/12% of the amount borrowed. Interest per month = 8/12 * 1/100 * 7,000 Interest over three months = 3 times above amount.

OpenStudy (anonymous):

how much will she pay back on the 20th?

OpenStudy (anonymous):

See my solution set up above and compute the values.

OpenStudy (anonymous):

im not getting it

OpenStudy (anonymous):

If interest of 8% is per year, per month it will be 8/12%. Interest is the extra amount you pay per month just for the fact that you borrowed someone else's money. So, for borrowing 7,000, you pay 8/12% of that amount each month extra. Over 4 months, you pay 4 times that amount extra.

OpenStudy (anonymous):

Interest per month = 8/12 % of 7,000 = 8/12 * 1/100 * 7,000 = 560/12 = 46.67. Over 4 months, you pay 4 * 46.67.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

TOY TOY IM DOING THE SAME PROBLEMS